OpenStudy (anonymous):

solve each system of equations: y=x-2,y=4x+1

OpenStudy (anonymous):

When you are solving a system of equations, you are finding the point of intersection

OpenStudy (anonymous):

use method of comparison thus we have x-2 = 4x +1 -3x=3 x=-1

OpenStudy (anonymous):

using x=-1 in y=x-2 we have y=-3 thus x=-1 and y=-3

OpenStudy (anonymous):

thank u i have another one

OpenStudy (anonymous):

y=3x+1,y=5x-3

OpenStudy (anonymous):

Are you solving algebraically or graphically?

OpenStudy (anonymous):

welcome use a similar method

OpenStudy (anonymous):

3x+1=5x-3 3x-5x=-3-1 -2x =-4 x=2 y=3(2) +1=7 so x=2 and y=7

OpenStudy (anonymous):

I'm using algebra to solve the linear equation